PROTON TUNNELING DYNAMICS OF PORPHYCENE AND 2,7,12,17-TETRA-t-BUTHYLPORPHYCENE UPON S_{1} ←S_{0} EXCITATION

Y. NOSENKO, J. JASNY, M. PIETRASZKIEWICZ, and \underline{A. MORDZINSKI}, Institute of Physical Chemistry Polish Academy of Sciences and Laser Center 44, Kasprzaka, 01-224 Warsaw, Poland.